OpenStudy (anonymous):

g(x) = x + 4 and h(x) = x^2 - 3 g[h(x)] = _____. A.x + 1 B.x 2 + 3 C.x 2 + 1 D.x 2 + 8x + 19

HanAkoSolo (jamierox4ev3r):

you have to combine the two functions

OpenStudy (solomonzelman):

If you were to be asked to find \(\large\color{slate}{ g(\color{red}{D}) }\) having the function: \(\large\color{slate}{ g(x) = x + 4 }\) you would do this: \(\large\color{slate}{ g(\color{red}{D}) =(\color{red}{D}) + 4 }\) Now, however, they are asking for \(\large\color{slate}{ g[\color{red}{h(x)}] }\) which just means that you are plugging the h(x) into the g(x) for x. knowing that \(\large\color{slate}{ h(x) = x^2 - 3 }\) (which they explicitly tell you) So lets plug in h(x) for x, into the function g(x). \(\large\color{slate}{ g[\color{red}{h(x)}] = (\color{red}{x^2 - 3}) + 4 }\)

OpenStudy (solomonzelman):

See what you are asked to do? Can you simplify this?

OpenStudy (anonymous):

x^2+1?

OpenStudy (solomonzelman):

yes, that is correct \(\large\color{slate}{ g[\color{red}{h(x)}] = x^2+1 }\)

OpenStudy (anonymous):

okay I get this now thank you

OpenStudy (solomonzelman):

Sure:) ... you welcome !